Week of 170918

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Maarten (ALICE), Gavin (Computing), Marian (Network), Borja (Monitoring), Julia (WLCG), Vincent (security), Andrea (MW)
  • remote: Andrej (ATLAS), Christoph (CMS), Zoltan (LHCb), Xin (BNL), Luca L (CNAF), Vincenzo (EGI), David (IN2P3), Sang (KISTI), Dmytro (NDGF), Onno (NL-T1), Elizabeth (OSG), Jose (PIC), Gareth (RAL), Di (TRIUMF)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Normal activities, 400k jobs on average
    • Problems
      • Small percent of transfers to CERN EOS were failing with checksum timeout error.
        Made FTS service to reduce number of threads leading to some backlog in transfers from BNL.
        Fixed at CERN by a change to settings on the GridFTP gateways.
      • FTS optimizer behaviour discussed

  • CMS reports ( raw view) -
    • Overall good CPU utilization last week: ~130k cores for production + ~50k for analysis
    • Still high transfer activity
      • Overall backlogs went down (partly due to LHC in Machine Development)
      • A few transfer issues being followed with various sites (nothing major as it seems)

  • ALICE -
    • Continued very high activity
    • CERN: LSF stopped scheduling jobs on Sat (GGUS:130599)
    • CERN: EOS-ALICE major upgrade this morning

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data processing and user analysis
    • Site Issues
      • T1: Failed transfers from IC to SARA (IPV6) (GGUS:129946); "Geant have confirmed that they are unable to ping mouse1.grid.sara.nl from geant-lhcone-gw.mx1.lon.uk.geant.net"

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF: NTR
  • EGI: NTR
  • FNAL:
  • IN2P3: Site in scheduled maintenance tomorrow.
  • JINR: NTR
  • KISTI: NTR
  • KIT:
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI:
  • OSG: Elizabeth mentioned the mail sent by Andrea S concerning the IPv6 deployment in WLCG.
    • Maarten: OSG will handle their T2 sites and report progress and issues through the IPv6 task force.
  • PIC: Scheduled downtime tomorrow for dCache update.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • ongoing move of 10k cores from LSF to Condor
  • CERN storage services:
    • EOSALICE upgraded to latest EOS release (codename Citrine) and XRootD 4.7)
  • CERN databases:
  • GGUS:
  • Monitoring:
    • Final reports for the August 2017 availability sent around
  • MW Officer: NTR
  • Networks: work in progress with GEANT for the network issue with Imperial College
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2017-09-18 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback